### v1

# Automating Knowledge Discovery from Scientific Literature via LLMs: A Dual-Agent Approach with Progressive Ontology Prompting

Abstract
To address the challenge of automating knowledge discovery from a vast volume of literature, in this paper, we introduce a novel framework based on large language models (LLMs) that combines a progressive ontology prompting (POP) algorithm with a dual-agent system, named LLM-Duo, designed to enhance the automation of knowledge extraction from scientific articles. The POP algorithm utilizes a prioritized breadth-first search (BFS) across a predefined ontology to generate structured prompt templates and action orders, thereby guiding LLMs to discover knowledge in an automatic manner. Additionally, our LLM-Duo employs two specialized LLM agents: an explorer and an evaluator. These two agents work collaboratively and adversarially to enhance the reliability of the discovery and annotation processes. Experiments demonstrate that our method outperforms advanced baselines, enabling more accurate and complete annotations. To validate the effectiveness of our method in real-world scenarios, we employ our method in a case study of speech-language intervention discovery. Our method identifies 2,421 interventions from 64,177 research articles in the speech-language therapy domain. We curate these findings into a publicly accessible intervention knowledge base that holds significant potential to benefit the speech-language therapy community.
footnotetext: Preprint. Under Review.
Introduction
With millions of research articles published annually, the overwhelming amount of existing scientific knowledge poses extreme challenges and opportunities for researchers to access knowledge through advanced analytical tools and interdisciplinary approaches. Knowledge discovery from scientific literature enables researchers to keep abreast of the latest developments in their domain and facilitate valuable insights that can significantly enhance their work (Usai et al. 2018; Wang et al. 2023). However, in such a vast ocean of data, only a very limited fraction of knowledge is collected and curated due to the low efficiency of the human review process. For example, in healthcare, evidence-based intervention refers to practices and treatments that are grounded in systematic research and have been proven effective through controlled studies (Rutten et al. 2021; Melnyk and Fineout-Overholt 2022). It emphasizes the use of evidence from well-designed and well-conducted research as the foundation for healthcare decision-making (Sackett 1997). One of the biggest challenges for healthcare providers is efficiently identifying relevant intervention evidence from research articles, thus underscoring the necessity of automating knowledge discovery.
Advancements in large language models (LLMs) present a significant opportunity for automating knowledge discovery in scientific literature. LLMs have been employed to categorize research papers, extract key findings, summarize complex studies, and streamline the review process effectively (Achiam et al. 2023; Guo et al. 2023). Some recent research (Li et al. 2023; Kim et al. 2024) utilize LLMs in conjunction with human annotators to reduce the human annotation burden. However, facing such a vast volume of domain knowledge, automating knowledge discovery is necessary to process and analyze information efficiently. For automating knowledge discovery from literature with LLMs, one big challenge is the limited context window length of LLMs (Wadhwa, Amir, and Wallace 2023). This constraint restricts the amount of input text that the models can process at one time, potentially leading to incomplete analysis and missed connections between data points spread across larger documents. To address this issue, the Retrievel-Augmented Generation (RAG) technique (Lewis et al. 2020; Gao et al. 2023) can be instrumental by combining a powerful retrieval component with a generative model, allowing the system to access a broader corpus of information beyond the immediate context window of a single model.
In this paper, we address the challenge of automating knowledge discovery through LLMs and formulate it as a problem of prompt design and scheduling based on a predefined ontology graph structure. To tackle this problem, we introduce a progressive ontology prompting (POP) algorithm that employs an outdegree-prioritized breadth-first search (BFS) across the ontology graph to create a series of prompt templates and action orders to guide LLMs. Additionally, to enhance the annotation quality, we further propose LLM-Duo, an interactive annotation framework for domain-specific knowledge discovery by leveraging the power of LLMs while addressing the limitations of LLMs. Specifically, this framework integrates two LLM agents working both collaboratively and adversarially to improve the discovery quality: 1) explorer, a chatbot based on the RAG technique, working in a zero-shot setting to generate annotation results while arguing with the evaluator to justify its answers; and 2) evaluator, an LLM tasked with evaluating annotation and providing feedback to assist explorer to refine its annotations.
To demonstrate the effectiveness of our approach in practice, we apply our method in a practical scenario, using speech-language intervention discovery as a case study. In our experiments, we compare our annotation framework to a number of advanced baselines including long context LLM (i.e., GPT-4-Turbo with 128k context window length), and RAG-based chatbot with prompting methods including Chain-of-Thought (CoT) (Wei et al. 2022) and SelfRefine (Madaan et al. 2024). The results show that our framework achieves more accurate and complete discoveries over baselines. The main contributions of our work are outlined as follows:
  • We formulate automated knowledge discovery with LLMs as a prompt design and scheduling problem, design a novel POP algorithm to transform a knowlege graph (KG) ontology into structured prompts and action orders, and achieve automatic knowledge discovery from literature.
  • We propose a novel annotation framework with two LLM agents that collaborate and compete to refine knowledge findings, which achieves superior performance over advanced baselines.
  • We apply our method to speech-language intervention discovery and successfully curate an intervention knowledge base that can serve as a valuable resource for the speech-language therapy community.
Related Work
Recent research demonstrates the impressive performance of LLMs in zero-shot prompting across various tasks such as natural language inference, question answering, summarization, named-entity recognition, and sentiment analysis (Agrawal et al. 2022; Kojima et al. 2022; Qin et al. 2023; Zhong et al. 2023b). Considering LLMs’ limitations (Abid, Farooqi, and Zou 2021; Bang et al. 2023), some work explores the possibility of LLMs and humans working collaboratively to annotate the same dataset. However, mining structured knowledge from long-range unstructured text is much more challenging for human annotators due to the vast volume of content, the inherent ambiguity of natural language, and the individual bias of human interpretation (Ye et al. 2022). These challenges are further compounded by the need for specialized expertise, labor-intensive manual annotation, and difficulties in maintaining consistency and scalability.
Traditional methods for knowledge graph construction (KGC) rely on NLP techniques, using pipelines for named entity recognition and relation extraction (Martins, Marinho, and Martins 2019; Wei et al. 2019; Zhong et al. 2023a; Laurenzi, Mathys, and Martin 2024). Recent advancements leverage LLMs to generate relational triplets in zero/few-shot settings. For example, ChatIE (Wei et al. 2023) uses multi-turn zero-shot question-answering for triplets extraction, achieving promising results. Some studies (Zhang and Soh 2024; Carta et al. 2023) further automate KGC by decomposing it into several phases and allowing LLMs to infer knowledge graph schemas without predefined ontologies. However, these methods are typically limited to processing short texts and have not been tested in practical scenarios. For long-form, domain-specific texts, expert-crafted ontologies can guide LLMs extract expected knowledge (Cauter and Yakovets 2024). Our approach introduces a novel method for constructing domain-specific knowledge graphs automatically using LLMs, which we detail in the following sections.
Preliminaries
Knowledge graph (KG) is a semantic network structured as an ontology, comprising concepts and their relationships with clear, interpretable structures at scale (Ehrlinger and Wöß 2016; Peng et al. 2023). For knowledge discovery from literature, LLMs can enhance this process by leveraging their capabilities to understand long-range text. This allows for transforming unstructured data into structured formats, and finally, populating the ontology to create KG.
Focusing on knowledge discovery from literature, we aim to gather domain-specific knowledge for researchers and practitioners across various fields. In our approach, the ontology is predefined by domain experts, which can be represented by a directed acyclic graph (DAG) $\mathcal{G}=(\mathcal{E},\mathcal{R},\mathcal{F})$. Here $\mathcal{E}$, $\mathcal{R}$, and $\mathcal{F}$ are sets of concepts, relationships, and semantic triples respectively. $\mathcal{F}$ is a collection of triples $(h,r,t)$ with a head concept $h\in\mathcal{E}$, a tail concept $t\in\mathcal{E}$, and a relation $r\in\mathcal{R}$ (Gruninger 1995). Named entity recognition (NER) and relation extraction (RE) are foundational elements for constructing knowledge graphs. To effectively instruct LLMs for knowledge discovery anchored to $\mathcal{G}$, the design of prompts and the sequencing of queries for both NER and RE tasks are essential for automating knowledge graph completion. We frame the problem of automated knowledge discovery via LLMs as one of prompt design and scheduling, described by the following equation:
$f(\mathcal{G}(\mathcal{E},\mathcal{R},\mathcal{F}))=\left\{(Prompt_{i},Order_{i})|i\in[1,N]\right\}$ (1)
where $f$ is a function that translates KG ontology into a set of prompt and action orders for the LLMs. A simple case of $f$ is directly injecting the whole ontology into a single prompt to instruct LLMs to generate KG annotation at once such as the annotation methods used in (Mihindukulasooriya et al. 2023; Kommineni, König-Ries, and Samuel 2024; Mihindukulasooriya et al. 2023).
Methodology
In this section, we first introduce the POP algorithm to convert a predefined knowledge ontology into a set of prompt templates and action orders, then propose an interactive annotation framework based on two LLM agents to enable more convincing and accurate annotation generations.
Ontology Prompting Algorithm
Figure 1: Illustration of prompt design and scheduling based on outdegree-prioritized BFS over a simple healthcare ontology for automating knowledge discovery.
Algorithm 1 Progressive Ontology Prompting
1:function OntologyTraversal($\mathcal{G}$, $k$)
2:     Initialize: $queue,Order\leftarrow[]$, $visited,Khops\leftarrow\{\}$
3:     Identify the source node $s$
4:     $enqueue(queue,(s,0))$, $visited[s]\leftarrow\text{true}$, $Khops[s]\leftarrow[]$
5:     while not $queue$.empty() do
6:         $(current,hop)\leftarrow dequeue(queue)$
7:         $Order$.append$(current)$
8:         Gather unvisited neighbors of $current$ into $neighbors$
9:         Sort $neighbors$ by out-to-in ratio in descending order
10:         for $neighbor$ in $neighbors$ do
11:              if not $visited[neighbor]$ then
12:                  $enqueue(queue,(neighbor,hop+1))$
13:                  $visited[neighbor]\leftarrow\text{true}$
14:                  $Khops[neighbor]\leftarrow[]$
15:              end if
16:              if $hop+1\leq k$ then
17:                  $Khops[neighbor].append(current)$
18:              end if
19:         end for
20:     end while
21:     return $Order$, $Khops$
22:end function
23:function Main($\mathcal{G}$, $k$, $T$)
24:     Initialize $annotations\leftarrow\{\}$
25:     $Order,Khops\leftarrow\textsc{OntologyTraversal}(\mathcal{G},k)$
26:     while not $Order$.empty() do
27:         $current\leftarrow Order$.dequeue(), $context\leftarrow Khops[current]$
28:         if not $context$.empty() then
29:              $discoveries\leftarrow\{annotations[t]\text{ for }t\text{ in }context\}$
30:              $prompt\leftarrow T(current,G_{sub}(context),discoveries)$
31:         else
32:              $prompt\leftarrow T(current)$
33:         end if
34:         $annotations[current]\leftarrow\textsc{LLM}(prompt)$
35:     end while
36:end function
In our work, we develop a progressive ontology prompting (POP) algorithm that employs a prioritized BFS on the ontology graph $\mathcal{G}(\mathcal{E},\mathcal{R},\mathcal{F})$ to generate a set of prompt templates and action orders for LLMs. As illustrated in Figure 1, in our method, the prompt design and scheduling follow a progressive manner. The annotation process begins by annotating the concept at a source node (i.e., nodes that only have outgoing edges) and then proceeds to its neighbors following the traversal order of a prioritized BFS. To allow BFS to quickly access a large portion of the graph, we modify BFS by sorting the neighbor nodes according to the out-to-in ratio $R(v)$, which is defined by:
$R(v)=\frac{|\left\{(h,r,t)\in\mathcal{F}\right\}|h=v|}{|\left\{(h,r,t)\in\mathcal{F}\right\}|t=v|}$ (2)
Therefore, the algorithm will choose the neighbor node with higher $R$ to visit at the next step. For example, in Figure 1, visiting the ‘Patient’ node before the ‘Disorder’ node can provide more information for annotating the ‘Disorder’ concept. For any concept node $v$, we use the visited nodes within its k-hop neighborhood as its context. The $Prompt_{v}$ for identifying concept $v$ is crafted based on its own concept type and knowledge discoveries within its context. The action order of this prompt $Order_{v}$ is determined by the order in which node $v$ is visited during the prioritized BFS traversal.
The pseudocode of the POP algorithm is shown in Algorithm 1. The algorithm first follows prioritized BFS traversal to capture the context and visit order of a specific concept node, then composes annotation prompts based on the local ontology structure and discoveries within its context as follows:
$Prompt(v)$$\leftarrow\left\{Prefix\left(N_{k-1}(u)\right)\oplus\right.$
$\left.Question\left((v,e,u)\mid(v,e,u)\in F\right)\mid u\in N_{1}(v)\right\}$ (3)
where $\oplus$ is the concatenation. We introduce two parameters in POP: context size $k$ and prompt template $T$. As the prompt examples shown in Figure 1, to annotate a specific concept, in prompt template $T$, we task LLMs to describe the knowledge discoveries within its context as a prefix to the annotation question associated with its concept type. Example prompts and corresponding outputs can be referred to in Appendix 3.
LLM-Duo Annotation Framework
Figure 2: Illustrative annotation process with LLM-Duo framework.
Figure 3: Two annotation examples of speech-language intervention discovery using LLM-Duo framework. These examples demonstrate the adversarial annotation process with two LLMs when separately emphasizing on rationality and completeness.
To guarantee the integrity and reliability of knowledge discoveries via LLMs, we propose LLM-Duo. The architecture of LLM-Duo is illustrated in Figure 2. It comprises two primary LLM components: the explorer and the evaluator. Specifically, the explorer is a chatbot developed with RAG and works for the intervention discovery and annotation by zero-shot question answering (QA). RAG is utilized to ground its generations within the given literature context. Even though RAG can reduce the hallucination of LLMs by providing references to guarantee the integrity and reliability of explorers’ generations, the explorer might still make mistakes. To ensure maximized accuracy and reliability of annotations, we introduce another LLM, called the evaluator, tasked with inspecting the explorer’s responses to enhance the annotation quality.
In our annotation framework, the interaction between explorer and evaluator is collaborative and adversarial. Given a research paper, LLM-Duo will be tasked with prompts following a specific action order obtained from the POP algorithm. Depending on the exact requirements of annotations for different knowledge domains, the inspection mechanism of the evaluator can be customized. In our test case of speech-language intervention discovery, we emphasize the factual correctness and completeness of the annotations for different concepts, which are also typical requirements for annotation quality in most scenarios. LLMs cannot directly assess the correctness of the answer since they do not have prior knowledge of the correct response. Inspired by the concept of LLMs self-correcting through reasoning, we convert the correctness inspection to the answer’s rationality assessment.
In an interaction cycle of LLM-Duo for annotating concepts emphasizing rationality, the answer along with its explanatory reason is presented to the evaluator. The evaluator scrutinizes the answer’s rationality and gives feedback to the explorer. Depending on the feedback, the explorer may either adjust its answer accordingly or, if disagreeing, present stronger evidence to support its original answer to challenge the evaluator’s assessments. For annotating concepts emphasizing completeness, in each round, the evaluator will extract the aspects covered in the explorer’s answer and integrate them with aspects of the last round to form a new aspect collection and greedily prompt the explorer to delve deeper and explore more beyond the new aspect collection. The iterative refinement process will continue until the annotation achieves consistency. The key idea behind our LLM-Duo framework is using the feedback loop to assist the explorer refine its annotations. As an example shown in Figure 3, by facilitating interactive loops between two LLM agents, LLM-Duo enables more convincing and accurate annotations.
Experiments
Implementation
In the LLM-Duo framework, the explorer is a chatbot built on LLM with RAG using Llamaindexhttps://www.llamaindex.ai. We use OpenAI ‘text-embedding-3-large’ as the embedding model and set the chunk size to 256 tokens with an overlapping size of 128. Particularly, we use ‘FastCoref’ (Otmazgin, Cattan, and Goldberg 2022) to process text chunks for coreference resolution before text embedding. Besides, we pass the document ID as metadata and apply a metadata filter on the chat engine to ensure the explorer only answers within the text of the specific document to be annotated. We set the retrieval to be on the top 8 text chunks based on similarity scores reranked with SentenceTransformerRerankhttps://docs.llamaindex.ai/en/stable/examples/node˙postprocessor/SentenceTransformerRerank employing the ‘cross-encoder/ms-marco-MiniLM-L-2-v2’https://huggingface.co/cross-encoder/ms-marco-MiniLM-L-2-v2 model in Llamaindex. The evaluator is an external LLM who does not share any document context with the explorer.
Case Description: Speech-language Intervention Discovery
Speech-language therapy provides interventions for individuals with speech-language deficits, enhancing their quality of life across various life stages. When choosing an intervention, evidence-based practice (EBP) is attractive as it integrates research evidence from literature into the decision-making process to ensure high-quality patient care (Law et al. 1996). Intervention research, especially studies that offer clear intervention frameworks and comprehensive case studies, are valuable references to guide EBP designs. Intervention discovery aims to extensively gather speech-language interventions from the literature corpus as references to facilitate EBP design. It involves identifying relevant studies and extracting essential features of interventions including target disorder, procedure, efficacy, case study, therapy activity, etc., which is extremely labor-intensive for human reviewers, highlighting the efficiency of automating knowledge discovery based on LLMs.
To verify the effectiveness of our method in a realistic scenario, we employ our framework in a speech-language intervention discovery setting. The ontology is shown in Figure 4. We list the concept and relation definitions of this ontology in Appendix 1.1. To enable a large-scale discovery, we cultivate a literature base including 64,177 papers within the domain of speech-language therapy. Specifically, we employ a set of keywords carefully selected from a glossary of commonly used terms in speech-language therapy to guide our literature search. The search keywords and literature base are shown in Appendix 1.2.
Figure 4: Ontology of speech-language intervention.
Baselines
The basic idea of LLM-Duo is to make LLM rectify its initial annotation results through external feedback from another LLM. Instead of setting up another LLM for evaluation, recent studies show that LLMs can perform self-correction to refine their responses by themselves (Liu et al. 2024; Li et al. 2024), and successful examples include Chain-of-Thought (CoT) (Wei et al. 2022) and SelfRefine (Madaan et al. 2024). We separately equip the chatbot based on RAG with these two prompting methods for annotation and denote them as CoT-RAG and SelfRefine-RAG. The full prompts for these two methods can be found in Appendix 2. Additionally, in LLM-Duo, a potential substitution of explorer is long-context LLM, which is capable of processing entire document tokens instead of chunking and retrieval with RAG. We denote LLM-Duo respectively using explorers built on RAG and long-context LLM as LLM-Duo-RAG and LLM-Duo-LongContext. Besides, we also compare with methods that directly input paper text to LLMs for zero-shot QA annotation without an evaluation feedback loop, including ShortContext LLM, LongContext LLM, OpenAI Assistant, and RAG.
(a) Illustration of participant annotation under different k values.
(b) Range of participant annotation accuracy using LLM-Duo-RAG with different LLMs at k=1,2,3.
(c) Token count distribution of retrieved-back chunks across varying similarity scores under participant annotation queries of k=1,2,3.
Figure 5: Overview of participant annotation and related metrics across different k values.
Evaluation
In our experiment of comparing with baseline approaches, we report six types of metrics: 1) Consistency Rounds (CR): the number of refine loops the method makes before achieving the annotation consistency; 2) Verbosity Index (VI): the number of aspects per 1k tokens of the annotations, which is an important metric for annotation emphasizing content completeness; 3) Enumeration Quantity (EQ): the number of items listed in the annotations (i.e., therapy activities, therapy goals in our test case.); 4) Faithfulness (Faith): the extent of the annotation faithful to the provided text, which is measured by FaithfulnessEvaluatorhttps://docs.llamaindex.ai/en/stable/examples/evaluation/faithfulness˙eval of Llamaindex. 5) Accuracy (ACC): the percent of correct annotations in all LLM-provided annotations. 6) Cover: the percent of correct LLM-provided annotations to the total mentioned concept entities in the provided text.
Results
In this section, we first provide a detailed evaluation of our progressive ontology prompting algorithm and the LLM-Duo annotation framework. Then, we showcase the results of speech-language intervention discoveries using our automated knowledge discovery framework, which further demonstrates the effectiveness of our framework in addressing practical scientific knowledge discovery tasks.
Context Size Analysis
In our progressive ontology prompting algorithm, when composing the annotation prompt for a concept node, the knowledge discoveries and the local ontology structure within the context (visited nodes within $k$-hop neighborhood) of a specific concept node serve as a prefix to provide conditions for annotation questions. The $k$ value determines the diversity and volume of information provided. To evaluate how the context size of the POP algorithm affects annotation quality, we conducted experiments by applying different $k$ values to generate prompts for LLM-Duo-RAG annotation. As shown in Figure 5(a), we selected an ontology substructure related to the ‘participant’ concept for the experiment, which was based on a random selection of 8 speech-language therapy papers.
Methods LLM IR ICA IKC
CR ACC Cover CR VI EQ Faith CR ACC
ShortContext GPT3.5-turbo - 36.9% 50% - 0.0249 5.46 0.9667 - 48.2%
OpenAI Assistant GPT4-turbo - 76.1% 69.0% - 0.0631 4.17 0.7857 - 53.3%
LongContext GPT4-turbo - 76.3% 57.1% - 0.0919 8.64 1.0 - 61.2%
LLM-Duo-LongContext GPT4-turbo 2.17 81.0% 68.7% 2.5 0.0926 8.68 0.8571 1.31 69.6%
RAG GPT3.5-turbo - 47.6% 50% - 0.0319 7.96 0.8550 - 48.7%
CoT-RAG GPT3.5-turbo 1.04 78.6% 81% 3.18 0.0771 10.37 0.7250 1.07 73.2%
SelfRefine-RAG GPT3.5-turbo 1.19 78.5% 54.4% 2.85 0.0694 7.17 0.8125 1.12 54.8%
LLM-Duo-RAG GPT3.5-turbo 1.84 100% 86.4% 2.58 0.1159 13.71 0.9285 1.46 85.6%
Llama3-instruct-70b 2.71 78.6% 55.6% 2.59 0.0748 9.79 0.8648 1.52 61.0%
Mistral-instruct-8x22b 2.30 81.9% 67.5% 2.16 0.0763 9.87 0.8875 1.46 67.2%
Table 1: Comparison with baseline methods using various base LLMs.
(a) Ablation studies of removing rerank (wo_rerank) and corefenence (wo_coref) modules in LLM-Duo-RAG.
(b) Number of discovered interventions for top 20 disorders.
(c) Number of discovered interventions for different age groups.
Figure 6: Ablation study and discovered speech-language intervention statistics.
The resulting annotation accuracy of the participant concept is shown in Figure 5(b). The results indicate that as the context hop $k$ increases, annotation accuracy improves significantly, suggesting that a larger context provides more informative prompts, thereby enhancing annotation quality. Moreover, GPT-4-turbo consistently outperforms GPT-3.5-turbo across all $k$ values, demonstrating that more advanced language models can further improve annotation accuracy. Besides, we inspect the text chunks retrieved back by different annotation queries based on various $k$ values. We report the range of similarity scores and token count distribution of retrieved-back chunks in Figure 5(c). The similarity score represents the semantically relevancy between retrieved texts to annotation queries. From the results, we can observe that when $k=1$, the retrieved text chunks generally exhibit low similarity with the query, and the token count tends to decrease as the similarity score increases, leading to poor annotation quality. In contrast, larger $k$ values, particularly $k=2$, result in more relevant content being retrieved. For $k=2$, the token count shows an increasing trend with higher similarity scores, indicating that more substantial and relevant content is being captured, which results in better annotation quality.
LLM-Duo with Baseline Annotation Methods
To evaluate the performance of our LLM-Duo annotation framework, we compare it with baseline methods in the context of speech-language intervention discovery. We randomly pick 8 papers from our speech-language literature base for annotation using various baselines. The comparison focuses on three key dimensions: 1) Intervention Recognition (IR), identifying intervention entities within the paper; 2) Intervention Aspect Summary (IAS), annotating an aspect (e.g., procedure, therapy activity, therapy goals) of an intervention approach, which requires capture and summarize all related information fragments from the paper; and 3) Intervention Knowledge Completion (IKC), linking interventions to theme classes (e.g., speech awareness, speech articulation, comprehension, foundation skills, etc.) and setting concept nodes (e.g., home, healthcare facilities, early childhood centers, schools, teletherapy, etc.). For IR and IKC tasks, we ask human annotators to complete annotation as golden results for comparison. For the IAS task, due to the individual bias of human interpretation, we only ask human annotators to tag the text fragments in the paper pertaining to a specific intervention aspect.
The experimental results are reported in Table 1. It should be noted that we implemented ‘ShortContext’ using Llama3-instruct-70b (FP16) and Mistral-instruct-8x22b models (INT8). However, directly prompting these models with full paper text fails to produce annotations in a zero-shot QA setting. Their generations do not align with the annotation questions. From the results in Table 1, we observe that LLM-Duo-RAG outperforms all baselines. Even though GPT4-turbo has a 128k context window length and is capable of generating annotations, its annotation coverage remains inadequate. Integrating it with the LLM-Duo framework can significantly improve both the accuracy and the comprehensiveness of the annotations. Additionally, compared with simple RAG, self-correct prompting methods such as CoT and SelfRefine can significantly enhance annotations, but their performance is still worse than LLM-Duo-RAG. Instead of utilizing costly GPT models, LLM-Duo-RAG, which employs open-source models including Llama3-instruct-70b and Mistral-instruct-8x22b, can achieve comparable annotation quality.
Ablation Study
In our framework, the chatbot explorer uses the RAG technique. We enhance generation quality with ’FastCoref’ for coreference resolution and rerank retrieved chunks by similarity score using the ’cross-encoder/ms-marco-MiniLM-L-2-v2 model’. This section presents ablation studies for both components. Since recognizing the intervention entity is the critical first step following the POP algorithm, we report the accuracy of intervention recognition in this study. As shown in Figure 6(a), the results demonstrate that removing these components significantly decreases annotation accuracy, showing the necessity of each module.
Speech-Language Intervention Discovery
Our knowledge discovery framework has identified 2,421 interventions supported by case studies from 64,177 research papers in the field of speech-language therapy. The statistics of discovered interventions are presented in Figure 6(b) and Figure 6(c). A complete intervention annotation example is available in Appendix 3. 19 clinicians and students reviewed our annotations through online Google forms. In checking the intervention annotation and reading the corresponding literature, they reported that 89% of our annotations are valid and complete. Based on the intervention ontology in Figure 4 and annotations, we have constructed an intervention knowledge graph that will be publicly accessible. This knowledge graph is expected to be a valuable resource for speech-language domain experts, supporting evidence-based clinical decision-making, question-answering and recommendation, ultimately enhancing healthcare outcomes.
Conclusion
In this paper, we developed a novel automatic knowledge discovery framework based on LLMs, which is characterized by a progressive ontology prompting algorithm and a dual-agent annotation system. The proposed method achieves superior performance compared with advanced baselines, enabling more accurate knowledge discovery. Applied to speech-language intervention discovery, this framework also curates a valuable, accessible intervention knowledge base for the speech-language therapy community.
Appendix A Appendix
1 Intervention Ontology and Literature
1.1 Intervention Ontology The intervention ontology is shown in Figure 4 of the main paper. In this section, we provide a detailed explanation of the concepts introduced in the ontology as follows:
  • Intervention represents a targeted treatment practice designed to enhance an individual’s communication skills.
  • Disorder represent the type of disorder that causes difficulties in an individual’s voice, speech, language, or swallowing functions.
  • Setting represents a specific environment where interventions are implemented. We identify six key settings: home, healthcare facilities (such as hospitals or rehabilitation centers), early childhood centers (like nurseries or daycare), schools, clinics and private practices, and teletherapy.
  • Theme represents the theme of the intervention. As shown in Table 2, we categorize interventions into 10 themes based on their characteristics and therapy goals.
  • Therapy Activity represents a task designed to address a particular speech or language challenge in an individual, such as using a minimal pairs activity to enhance phonological awareness.
  • Therapy Goal represents a specific area that the intervention is designed to enhance.
  • Procedure represents a comprehensive description of how the intervention is carried out.
  • Efficacy represents the conclusion about the effectiveness of intervention.
  • Frequency/Dosage/Duration represents the frequency/dosage/ duration of the intervention practiced in the case study that demonstrates its efficacy.
  • Case Study represents a detailed examination of the intervention on a particular individual or group with communication disorders. The purpose of a case study is to provide a deep understanding of the patient’s unique needs and assess the intervention effectiveness.
  • Participant represents the individuals or populations that are involved in the case study of intervention.
  • Age represent the age of experiment participant or claimed target population of the intervention. The age is quantified with a granularity of half a year. We additionally convert age to age groups including “newborn, infants, toddlers, children, preschoolers, school-age children, older children, youth, teens, adolescents, adult, young adult, middle aged, aged, senior”. Each specific age may be associated with multiple age groups. For instance, an individual aged 13 years could be categorized to the ‘teens,’ ‘adolescents,’ and ‘children’ age groups.
  • Language represent the speaking language of experiment participant in the case study of the intervention.
Theme Definition
speech awareness work that involves recognizing and understanding speech sounds. It includes phonological awareness (identifying and manipulating sounds), auditory discrimination (distinguishing sounds), and sound identification (recognizing the meaning of sounds).
speech articulation work that refers to the physical production of speech sounds using the mouth, lips, tongue, and respiratory system. It focuses on the clarity and accuracy of pronouncing phonemes and forming them into words.
comprehension work that aims to improve the understanding of (receptive) language.
expressive language work that aims to improve the children’s expressive language, in quantity, vocabulary or structure.
self-monitoring work designed to help the patient’s awareness of their speech and language difficulties and how they might be able to overcome them.
generalisation work to help make speech and language or therapy gains transferable to other situations and environments.
foundation skills work to practise and improve a range of early skills, many of which might be considered foundations for speech and language development.
functional communication work focusing on those aspects of communication that help the child’s involvement and participation in life situations; this might be functional language, signing or the use of symbols.
adult understanding and empowerment work that helps parents to understand the nature of their child’s speech and language difficulty, what helps to improve it and why.
adult-child interaction work on the interaction between the parent/adult and the child. All of the changes to adult/parent-child interactions were emphasised in terms of those that encourage speech and language development.
Table 2: Definition of intervention themes.
Source Number of Literature
Wiley Online Library 9,511
NIH PubMed 30,928
SpringerLink 6,549
GoogleScholar 8,662
ASHA 5,641
EBSCOHost 2,886
Table 3: Literature base for speech-language intervention discovery.
1.2 Speech-language Therapy Literature Collection We cultivate a literature corpus within the domain of speech-language therapy as shown in Table 3 to facilitate the discovery of intervention methods. To conduct our literature search, we use a collection of carefully selected keywords drawn from a glossary of commonly used terms in speech-language therapy. These keywords include:“speech language therapy, speech language disorder, speech sound disorder, articulation disorder, speech intervention, language intervention, auditory discrimination, auditory processing disorder, phonological awareness, phonological processes, auditory perception, babbling, motor speech disorder, morpheme, phonology, prosody, stuttering, language impairment, speech language pathologist, speech and language therapist, babbling, expressive language delay, cleft speech disorder, autism spectrum disorder, developmental phonological disorder, developmental stuttering, phonological impairment, developmental dysarthria, down syndrome, swallowing disorder, communication impairment, articulation impairment, dyslexia, apraxia, dysarthria, dysphagia, communication disorder, expressive language disorder, dyspraxia, aphasia, augmentative and alternative communication, central auditory processing disorder, cleft lip and palate, down syndrome, fluency disorders, hearing loss, orofacial myofunctional disorders, spoken language disorders, written language disorders, acquired brain injury, apraxia of speech, auditory comprehension, literacy impairments, voice difficulties, language-based learning disabilities.”
2 Prompting Methods
2.1 POP Prompting
In the POP algorithm, the annotation prompt for a specific concept node is generated from a template $T$, which is constructed using the node’s context (the k-hop visited neighborhood) and the knowledge discoveries within that context. We carefully design the prompt to task LLM generate $T$ as follows:
The following triplets outline an annotation ontology:{(Intervention, StudiedIn, Case Study), (Intervention, Include, Participant), (Case Study, UsedWith, Frequency)}. All concept nodes have been annotated except from {#Frequency}. Your task is to create all possible annotation prompt templates for the {#Frequency}, leveraging the ontology structure.
Example:
Ontology: [(Intervention, StudiedIn, Case Study), (Intervention, TargetAt, Disorder), (Case Study, Include, Participant), (Participant, Has, Disorder)].
Annotate: Participant
Prompts:
T1: Who is the participant in {#case study}?
T2: {#Intervention} is studied in {#case study}. Who is the participant in {#case study}?
T3: {#Intervention}, targeting {#Disorder}, is studied in {#case study}. Who is the participant in {#case study}?
Some example output of GPT3.5 for the above prompt is:
T1: What is the frequency used in {#case study}?
T2: In the context of {#intervention} studied in {#case study}, what is the frequency of the intervention?
T3: {#Intervention}, which includes {#participant}, is studied in {#case study}. What frequency is used in this study?
Since template $T$ is composed of concept nodes visited prior to the current node, the final annotation prompt for the current node is derived by populating the knowledge discoveries associated with these concept nodes into $T$.
2.2 CoT and SelfRefine Prompting
Instead of using an external LLM to provide evaluation feedback to explorer, we use CoT and SelfRefine prompting techniques as baselines to task explorer refine annotations independently. The prompts of CoT and SelfRefine are as follows:
Background: Your last answer to my question {#init annotation question} is: {#last annotation}.
*CoT:
Prompt_CoT1: "{#Background} Make a plan to correctly answer my question again."
Prompt_CoT2: "{#Background} Make a plan to answer my question again with more comprehensive results."
*SelfRefine:
Prompt_SelfRefine_Feedback1: "{#Background} Reflect your answer. Analyze the correctness of the information provided. Provide critque to help improve the answer. Your feedback:"
Prompt_SelfRefine_Refine1: "{#Background} Critics: {#feedback}. Based on your last answer and its critics, revise your answer to my question. Your answer:"
Prompt_SelfRefine_Feedback2: "{#Background} Reflect your answer. Analyze the included aspects in your answer. Provide critque to help make the response more comprehensive. Your feedback:"
Prompt_SelfRefine_Refine2: "{#Background} Critics: {#feedback}. Based on your last answer and its critics, revise your answer to my question. Your answer:"
3 Annotation Example
Below is an example of an intervention annotation based on the study titled ’Intensive Treatment of Dysarthria Secondary to Stroke.’
"Intervention": "Lee Silverman Voice Treatment (LSVT LOUD)",
"Disorder": [
"stroke",
"ataxia",
"multiple sclerosis",
"traumatic brain injury (TBI)"
"Procedure": "The therapy process of conducting the Lee Silverman Voice Treatment (LSVT LOUD) intervention involves intensive high-effort exercises aimed at increasing vocal loudness to a level within normal limits using healthy and efficient voice techniques. The treatment protocol includes sessions four times a week for 4 weeks, totaling 16 individual one-hour sessions. Each session consists of tasks such as maximal sustained vowel phonation, pitch range exercises, and reading functional phrases at individual target loudness levels. The second half of each session progresses to functional speech tasks, moving from words and phrases to conversation over the course of the 16 sessions. Additionally, participants are assigned daily homework to practice using normal loudness and facilitate generalization of normal loudness outside the treatment room.",
"Frequency": "four times a week",
"Dosage": "one-hour session.",
"Duration": "4 weeks, totaling 16 individual one-hour sessions.",
"Efficacy": "The outcome of the Lee Silverman Voice Treatment (LSVT LOUD) intervention in this study showed positive changes in acoustic variables of speech for all participants with dysarthria secondary to stroke. There were statistically significant increases in vocal dB SPL for sustained vowel phonation and speech tasks, indicating improvements in loudness levels and phonatory stability. Additionally, post-treatment speech samples were rated as having better voice quality and sounding more natural, suggesting an amelioration of dysarthria characteristics. Participants also reported increased confidence in their speech during post-treatment interviews.",
"Therapy Goal": [
"Increase vocal loudness to a level within normal limits",
"Use healthy and efficient voice techniques",
"Progress from words and phrases to conversation over 16 sessions",
"Facilitate generalization of normal loudness outside the treatment room"
"Participant": "Four participants (P01 to P04).",
"Age": "Participants in the study ranged in age from 50 to 74 years.",
"Language": "assume to be English.",
"Case Study": "Case studies and experiments regarding the Lee Silverman Voice Treatment (LSVT LOUD) intervention in the paper include studies on people diagnosed with stroke, ataxia, multiple sclerosis, and traumatic brain injury (TBI). These studies have shown improvement in articulatory features as well as loudness.",
"Therapy Activity": [
"Maximal sustained vowel phonation",
"Pitch range exercises",
"Reading 10 functional phrases at individual target loudness levels",
"Functional speech tasks progressing from words and phrases to conversation"
"Setting": "home",
"Theme": "speech articulation"
4 Speech-Language Intervention Knowledge Graph
Below is the speech-language intervention knowledge graph constructed using our knowledge discovery framework. We plan to make this knowledge base publicly accessible after our paper is published.
Figure 7: Graph database overview.
Figure 8: Partial views of Intervention-Disorder in the intervention KG.
Figure 9: Partial views of Intervention-TherapyActivity in intervention KG.